OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

What do two lines that are parallel have in common? Since an equation of a line only has a slope and a y-intercept, what do two parallel lines have in common?

OpenStudy (anonymous):

i havent been in school for a long time and i do not remember

OpenStudy (kainui):

You might want to check this site out, they have a lot of awesome videos that explain things. This might be a good video to start with but you might be able to find better. https://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/slope-and-intercepts/v/slope-of-a-line

OpenStudy (anonymous):

y=mx+b to be parallel, slopes must be equal rearrange 4x-3y=-12 3y=4x+12 y=4/3x +4 slope (m) is 4/3 sub in points it has to pass through (x=-3, y=4) (4)=4/3(-3)+b y intercept (b) is 8 Equation: y=4/3x+8